2. Suppose that a large lot of lambs contains 17% defectives. Find the probability that a random sample of 8 fuses contains: a) Exactly two defective, b) At most two defective, b) At least one defective.
Answer #1

Solution:

We are given that a large lot of lambs contains 17% defectives. Therefore, we have:

p=0.17

Let x be the number of defectives

We know that a random variable x follows a binomial distribution, with parameters n=8,p=0.17

a) Exactly two defectives.

Answer:  

P(z = 2-C)0.172(1-0.17)8-2 P(z = 2) = 2

8! 0.1 0.17)2

28 × 0.0289 × 0.326940373

0.2646

Therefore, the probability of exactly two defectives is 0.2646.

b) At least one defective.

Answer:  

P(xgeq1)=1-P(x<1)

= 1-P(z = 0)

0.170(1-. 0.17)8-0 0

1-0.2252

=0.7748

Therefore, the probability of at least one defective is 0.7748.

c) At most two defective

Answer:

P(x leq 2)=P(x=0)+P(x=1)+P(x=2)

=inom{8}{0}0.17^{0}(1-0.17)^{8-0}+inom{8}{1}0.17^{1}(1-0.17)^{8-1}+inom{8}{2}0.17^{2}(1-0.17)^{8-2}

=0.225229+0.369050+0.264560

0,8588

Therefore, the probability of at most two defectives is 0.8588